# A Stochastic Block Hypergraph model

### v1

@@ IV Discussion
We proposed here a simple generalization of the stochastic block model for hypergraphs. Other models were proposed recently, but here we focus on the hyperedge formation process. When a new node is considered, it connects to existing hyperedges according to some probability. This probability depends on the structure of the hyperedge and reflects the formation process of a group of nodes. This is reminiscent of group formation in various systems: when a group exists, what is the probability that a new individual will join the group? In the framework proposed here, such a process can be explicitly integrated in the model. In the simple cases considered here, our results suggest that the degree and size distributions are approximately binomials with effective parameters that depend on the connection probability (in our case, they depend on $q/p$). We also found that the strategy has indeed an impact on the composition of hyperedges, and our results suggest that the dependence on the composition is a critical parameter in the evolution of hypergraphs. Also, the strategy has a strong impact on the heterogeneity among hyperedges and we even observe that the maximal heterogeneity could be obtained for an intermediate value of $q/p$ (and not necessarily for $q=p$). Our results also suggest that when the strategy depends on the composition of the hyperedge, the resulting hypergraph has hyperedges with a small diversity (essentially made of nodes from the same community). This is in contrast when the formation process is independent from the hyperedge structure, that leads to hyperedges comprising a larger diversity of communities.
Important advantages of this model are its small number of parameters and that it is easy to implement. For these reasons, we believe that model could be used as a benchmark for community detection methods for hypergraphs, but could be also useful to test various dynamics on hypergraphs.

### v2

@@ IV Discussion
We proposed here a simple generalization of the stochastic block model for hypergraphs. Other models were proposed recently, but here we focus on the hyperedge formation process. When a new node is considered, it connects to existing hyperedges according to some probability. This probability depends on the structure of the hyperedge and reflects the formation process of a group of nodes. This is reminiscent of group formation in various systems: when a group exists, what is the probability that a new individual will join the group? In the framework proposed here, such a process can be explicitly integrated into the model. In the simple cases we considered, our results suggest that the degree and size distributions are approximately binomials with effective parameters that depend on the connection probability (in our case, they depend on $q/p$). We also found that the strategy has indeed an impact on the composition of hyperedges, and our results suggest that the dependence on the composition is a critical parameter in the evolution of hypergraphs. Also, the strategy has a strong impact on the heterogeneity among hyperedges and we even observe that the maximal heterogeneity could be obtained for an intermediate value of $q/p$ (and not necessarily for $q=p$). Our results also suggest that when the strategy depends on the composition of the hyperedge, the resulting hypergraph has hyperedges with a small diversity (essentially made of nodes from the same community). This is in contrast when the formation process is independent from the hyperedge structure, that leads to hyperedges comprising a larger diversity of communities.
Important advantages of this model are its small number of parameters and that it is easy to implement. It can be used to model phenomena where the formation process is known or can be approximated and where we expect degree and hyperedge size distributions that follow binomial laws, such as in [53]. It can also be used as a null model or as a benchmark for several tasks, such as the detection and visualization of communities in hypergraphs.